The temperature coefficient of resistivity of a material is 0.0049 (C°)-1. By how much ΔT does the temperature of a wire made from this material have to change, in order that the resistance of the wire doubles?
Answer #1

= Temperature coefficient of resistivity = 0.0049 C-1

T = Change in temperature = ?

Ri = initial value of resistance = R

Rf = final value of resistance = 2R

Final resistance is given as

Rf = Ri (1 + T)

2R = R (1 + T)

2 = (1 + T)

2 = (1 + (0.0049)T)

T = 204.08 C°